In a two- slit experiment, monochromatic coherent light of wavelength 600 nm passes through a
pair of slits separated by 2.20 × 10 - 5 m. At what angle away from the centerline does the second
LIGHT fringe occur?
Answer #1


In double slit


path difference = d*sintheta = m*lambda

for second bright m = 2


d = slit seperation = 2.2*10^5 m

lambda = wavelength = 600*10^-9 m


2.2*10^-5*sintheta = 2*600*10^-9

theta = 3.13 degrees <<<=======ANSWER